OpenStudy (amtran_bus):

Find the equation of the line with the property: It has X-int of 4 and Y-int of 5

OpenStudy (campbell_st):

well you have 2 points (4, 0) and (0, 5) you can find the slope and then the equation of the line

hartnn (hartnn):

\(line : \\ \dfrac{x}{x-intercept}+\dfrac{y}{y-intercept}=1\)

hartnn (hartnn):

^thats the intercept form of line

OpenStudy (amtran_bus):

Thanks, so the answer is y=-5/4 +5?

OpenStudy (amtran_bus):

that is, -5/4 x

hartnn (hartnn):

yes, correct :)

OpenStudy (amtran_bus):

Thanks!